Question 34
ProbabilityA box contains 14 white balls and 6 black balls. Find the probability of first drawing a black ball and then a white ball without replacement.
- A.0.21
- B.0.22
- C.0.30
- D.0.70
Show the answer and explanation
Answer: B (0.22)
Explanation
6/20 × 14/19 = 84/380 ≈ 0.22.